A study was conducted concerning the effects of varying lengths of unemployment on ability to concentrate. Data represent participants' proofreading scores (number of errors found). Finding fewer errors on the proofreading task is a behavioral indicator of stress. Length of unemployment was regarded by the researchers as an indicator of exposure to chronic stress. Fifteen participants were included in the study and followed across three periods: employment (control), short-term unemployment and long-term unemployment. (20%) Data were collected and are shown below: Length of Unemployment

Control: 8 10 9 8 10. Short: 0 1 2 7 5 Long: 6 7 7 6 9

Please analyze the data fully (including appropriate follow-up tests) and provide a brief written summary of the results.
Answer #1

As there are more than 2 lengths of unemployement, we shall use an ANOVA to test the mean number of errors across the lengths of unemployment.

Total
Control 8 10 9 8 10 45
Short 0 1 2 7 5 15
Long 6 7 7 6 9 35
Total 95

Null hypothesis: The avaerage number of errors made in proof reading is the same across three employment levels.

Alternate Hypothesis: The average number of errors vary for at least two periods.

Level of significance: g= 0,05

Test Statistic: We shall use the F-test by constructing ANOVA table:

1. Calculate the correction factor: CF=\frac{\left (\sum_{i=1}^{3} \sum_{j=1}^{5}y_{ij}\right )^2}{n}

CF=\frac{95^2}{15}=\frac{9025}{15}=601.6667

2. Calculate the total sum of squares: TSS=\sum_{i=1}^{3} \sum_{j=1}^{5}y_{ij}^{2}-CF

We need the squares of individual values and total them:

  

Control 64 100 81 64 100 409
Short 0 1 4 49 25 79
Long 36 49 49 36 81 251
Total 739

TSS=739-601.6667

TSS=137.3333

3. The sum of squares due to periods: SSP

SSP=\sum_{i=1}^{3}\frac{y_{i\cdot }^{2}}{5}-CF

  SSP=\frac{45^2+15^2+35^2}{5}-601.6667

SSP=\frac{2025+225+1225}{5}-601.6667

SSP=\frac{3475}{5}-601.6667

SSP=695-601.6667

SSP=93.3333

4. Error sum of squares ESS=TSS-SSP

ESS=137.3333-93.3333

SSP=44

We shall concstruct the ANOVA table:

Source of Variation SS df MS F P-value F crit
Between Groups 93.3333 2 46.6667 12.7273 0.0011 3.8853
Within Groups 44.0000 12 3.6667
Total 137.3333 14

From the ANOVA table, since the p-value for teh periods is p=0.0011<0.05, we reject the null hypothesis. Hence, we conclude that the average number of errros differes between at least two periods of employment.  

We shall do a post-hoc analysis of the means to find out which pairs are differing significantly.

We shall calculate the Least significant difference using the formula

LSD=t_{1-\alpha /2,12}*\sqrt{MS_{errror}*\left ( \frac{1}{5}+\frac{1}{5} \right )}

LSD=2.1788*\sqrt{3.6667*0.4}

LSD=2.1788*\sqrt{1.4667}

LSD=2.1788*1.2111

LSD=2.6387

We shall form the differenc table as below:

Groups Average Comparison Absolute difference LSD Significant
Control 9 Control Vs. Short 6 2.6387 Yes
Short 3 Control Vs. Long 2 2.6387 No
Long 7 Short Vs. Long 4 2.6387 Yes

We can see from the above that the difference between Control and Short and Short Vs. Control are statistically different.